Fermi National Laboratory
 

Accelerator Update for December 15, 2006 - December 18, 2006

 

Friday December 15

The day shift began with Operations monitoring store 5127 (74.64E30) and a Recycler stash (177.1E10), with the Antiproton Source (Pbar) stacking (27.0E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 9:13 AM, Operations reported that the electron cooling system had tripped off due to a LCW leak. Operations had prepared the area for access and experts were already working on the system.

Operations prepared the Recycler and Pbar to stash antiprotons at 10:32 AM.

A NuMI kicker expert reported that the kicker temperature problem was due to a clogged water filter. The filter was changed at 10:55 AM.

Pbar resumed stacking at 11:47 AM.

A Switchyard (SY) expert began establishing beam to the SY absorber at 12:08 PM.

Recycler experts completed their work on the electron cooling LCW system at 2:45 PM.

MiniBooNE experts reported at 2:56 PM that they would resume normal horn operations later this afternoon. They had been running with their horn off for a week.

The Recycler completed a stash at 3:59 PM.

Operations received permission to send beam to the T959 experiment in MCenter at 4:31 PM. A Meson expert established beam to the M01 absorber at 5:43 PM.

The Recycler completed a stash at 6:36 PM.

The Meson expert established beam to the Meson T959 experiment at 8:28 PM.

The Recycler completed a stash at 11:51 PM.

Saturday December 16

The midnight shift began with Operations monitoring store 5127 and a stash (373.5E10), with Pbar stacking (27.5E10), and with T959, NuMI, and MiniBooNE taking beam.

Operations terminated store 5127 at 12:05 AM.

Operations put the TeV into shot setup at 12:41 AM.

Operations established store 5128 at 3:32 AM with an initial luminosity of 229.0E30. This store was Fermilab’s second highest luminosity.

The Recycler completed a stash at 4:14 AM.

The Recycler completed a stash at 6:44 AM.

At 7:46 AM, Operations reported that the I- Source had tripped off numerous times during the midnight shift.

At 7:54 AM, the NuMI horn tripped off on a ground fault indication. It reset okay.

At 8:41 AM, Operations disabled the T959 beam for a scheduled access into M02.

The Recycler completed a stash at 9:37 AM.

Operations reestablished beam to T959 at 9:57 AM.

At 10:25 AM, Operations disabled T959 beam for an access into M02.

The Recycler completed a stash at 12:25 PM.

Operations reestablished beam to T959 at 12:34 PM.

At 1:42 PM, Operations disabled T959 beam for an access into M02.

The Recycler completed a stash at 2:41 PM.

Operations reestablished beam to T959 at 2:51 PM.

At 3:18 PM, Operations reported that the I- Source had tripped off numerous times during the day shift.

The Recycler completed a stash at 5:22 PM.

At 7:44 PM, Operations disabled T959 beam for an access into M02.

The Recycler completed a stash around 8 PM.

Operations reestablished beam to T959 around 8:15 PM.

At 8:40 PM, store 5128 aborted. There was no quench. The Cryo System Run Coordinator assisted with the investigation.

At 9:13 PM, the MiniBooNE horn tripped off. It’s power supply failed. An expert came in and replaced the supply.

The Cryo Run Coordinator pointed out at 9:27 PM that there was a fluctuation in lead voltages and temperatures on a sector BA magnet (LDV3). Operators will change out a VFC card and then put the TeV through a dry squeeze.

At 10:21 PM, after operators had ramped down the Tevatron (TeV), the SY beam positions moved, taking beam away from T959. Operations contacted SY experts.

The Recycler completed a stash at 11:30 PM.

Sunday December 17

The midnight shift began with Operations monitoring the stash, and with T959, MiniBooNE, and NuMI taking beam.

Operations reported at 1:13 AM that they had changed the VFC card, conducted a dry squeeze and got the same lead trip. Operators swapped out the chassis and got the same trip. They swapped cards and got the same trip. Operators called in a controls expert.

The Recycler completed a stash around 1:38 AM.

Operations began preparing the TeV for an access at 2:14 AM. The plot for BA lead 2 showed the signs of a bad connector.

Cryo system technicians accessed TeV sector BA and reported at 5:18 AM that they had found an ice ball interfering with magnet connector. They melted the ice ball and repaired the connector.

Operations put the TeV through a dry squeeze at 6:04 AM.

The Recycler completed a stash at 6:48 AM.

Operations put the TeV into shot setup at 7:22 AM.

At 9:02 AM, Operations reported that they had lost two transfers of antiprotons.

Operations established store 5134 at 9:31 AM with an initial luminosity of 213.56E30. This was a 36x32 store (36 bunches of protons and 32 bunches of antiprotons).

At 8:11 PM, Operations reported that the I- Source had tripped off numerous times (106 times) today.

At 8:43 PM, a Booster RF station (BRF6) tripped off and wouldn’t reset. Operators bypassed the station.

Operations reported that this was the sixth and last stash for the day at 10:49 PM. T959 had their beam disabled and accessed M02 during each of these stashes and then came back up after the stash was completed.

Monday December 18

The midnight shift began with Operations monitoring store 5134 and the stash, and with T959, MiniBooNE, and NuMI taking beam.

The T959 experimenters reported at 1:13 AM that their experiment had completed its run. They will pack up their equipment in the morning. Operations disabled the beam line.

The Recycler completed its fourth and last shot for the midnight shift around 8 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The T959 experimenters completed their run and are removing their equipment.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/18/2006    email Fermilab